OpenStudy (anonymous):

(A) Find the parametric equations for the line through the point P = (-2, -1, 0) that is perpendicular to the plane −5x − 5y + 5z = 1. Use "t" as your variable, t = 0 should correspond to P, and the velocity vector of the line should be the same as the normal vector to the plane found directly from its equation. x = ??? y = ??? z = ??? (B) At what point Q does this line intersect the yz-plane? Q = ( ? , ? , ? )

OpenStudy (anonymous):

either this one is very simple like this (-2,-1,0)+t(-5,-5,5) BECAUSE N=(-5,-5,5) thats perpendicular to the plane

OpenStudy (anonymous):

Find the intersection of the line Q with the coordinate planes. For the y-z plane, we simply set x = 0 = -2 - 5tyz to find the value of the parameter t for the y-z plane: tyz = -2/5. Now we use this value of t in the equations for y and z: Yyz = -1 - 5tyz = -1 - 5(-2/5) = 1 Zyz = 5*tyz = 5(-2/5) = -2 In other words, Q intersects the y - z coordinate plane at (0,1,-2)

OpenStudy (anonymous):

the second part is right but how about A ?! i mean what is the X, Y and Z ?!

OpenStudy (anonymous):

X=-2-5t Y=-1-5t Z=5t or (x,y,z)=(-1,-1,0)+t(-5,-5,5)

OpenStudy (anonymous):

thanks !